Circular No. 7729

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001es IN ANONYMOUS GALAXY W. D. Li, University of California at Berkeley, reports the discovery

by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.0) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Oct. 7.4 and 8.4 UT. SN 2001es is located at α = 2^{h}02^{m}08^{s}.27, δ = +19^{o}05′27′′.5 (equinox 2000.0), which is 0′′.8 west and 2′′.5 north of the nucleus of a galaxy that is itself located ∼ 150′′ west and 84′′ north of the nucleus of UGC 1518. A KAIT image of the same field taken on Sept. 27.4 showed nothing at this position (limiting mag ∼ 19.0).

NOVA IN M31 Li also reports the discovery of an apparent nova in M31 on unfiltered

KAIT images taken on Oct. 5.4 (mag ∼ 17.5) and 6.4 UT (mag ∼ 16.9). The nova is located at α = 0^{h}43^{m}03^{s}.30, δ = +41^{o}12′11′′.3 (equinox 2000.0), which is ∼ 213′′ east and 238′′ south of the nucleus of M31 (= NGC 224). A KAIT image taken on Oct. 1.4 showed nothing at this position (limiting mag ∼ 19.0).

CYGNUS X-1 G. Pooley, Cavendish Laboratory, University of Cambridge, communi-

cates: “Recent observations of Cygnus X-1 with the Ryle Telescope show unusually low radio emission, with a mean flux density of ∼ 0.95 mJy at 15 GHz over the period Oct. 1–9. For the past five years, the source has usually been in a low/hard x-ray state, with radio emission always present at the level 5–40 mJy. The current lack of radio emission corresponds to a transition to the high/soft x-ray state, as seen in the publicly available RXTE monitoring data.”

M. P. Rupen, A. J. Mioduszewski, and V. Dhawan, National Radio Astronomy Observatory, further report: “Observations with the Very Large Array on Oct. 3 confirm the Ryle results noted above, putting even lower limits on the radio flux density, with no detections to rms noise levels of 0.58, 0.17, 0.12, and 0.34 mJy/beam at 1.425, 4.860, 8.460, and 14.939 GHz, respectively. Further observations of this unusual radio/x-ray state at other wavelengths are strongly encouraged.”

2001 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green